OpenStudy (anonymous):

Complete the table to determine the balance A for $14,000 invested at rate r = 4.5% for t years, compounded continuously. (Round your answers to two decimal places.) T-10,20,30,40,50

OpenStudy (callisto):

Use FV = PV (1+i)^n FV = amount you need to find PV = amount you invested i= interest rate n = no f years That will do if your case is just invest for one time only, not continuous investment

hero (hero):

Hint: Continuous Compound Interest Formula: \[A = P e^{rt}\]

OpenStudy (anonymous):

10=4.89 20=1.71 30=5.96 40=2.09 50=7.28 A=Pe^(r(t)) first find the rate and time then for example 4.5(10)=45 then 14000e^(45)=4.89 so on and so on

hero (hero):

Start with the first time: P = 14,000 r = .045 t = 10 \[A = 14000 e^{.045 \times 10}\] Just to be clear r = .045 not 4.5

OpenStudy (anonymous):

good catch hero

OpenStudy (anonymous):

OP follow the concept hero posted and that should give you the correct answers